# HoneySat: A Network-based Satellite Honeypot Framework

### v2

@@ VII Discussion and Future Work
Limitations. Currently, the functionality of some subsystems within HoneySat’s Satellite Simulator are constrained by the quality of the data provided. For example, the resolution of Earth’s images generated by our camera payload depends on the resolution of its source data (USGS [77]). Consequently, creating a honeypot for a satellite with a high-resolution camera payload would not be feasible without addressing the underlying issues of data source quality.
@@ VIII Conclusion
Although we have yet to witness a Stuxnet-like cyberattack on space systems, security researchers need to develop effective countermeasures to secure satellites. In this paper, we introduced HoneySat, the first satellite honeypot that realistically simulates an entire satellite mission and provided strong evidence that shows that our honeypot can obtain real-world interaction data and can be extended and customized with real-world satellite software and hardware. Finally, we hope that security researchers use HoneySat’s open-source implementation as a foundation not only for satellite honeypot deployments but also for simulation, education, and training applications.
@@ IX Ethical Considerations
In this paper, we consider the ethical consequences and possible negative outcomes of two main elements of our research. The satellite operator user study and the multiple deployments of our satellite honeypot exposed to the Internet. We now discuss the stakeholders, potential risks, and how we mitigated those risks for each of these elements.
@@ IX-A Satellite Operators User Study
Prior to conducting any research, the study plan, survey protocol, and associated materials were submitted for review by our Institutional Review Board (IRB), which evaluated the study’s potential risks to participants. The protocol received ”exempt status” from the IRB, indicating that the research involves no more than minimal risk. This status reflects our careful evaluation of possible harms, which we discuss further below.
Stakeholder Identification. The primary stakeholders in the survey are survey participants. Our study aims to gather voluntary responses regarding the fidelity and realism of our honeypot system. We carefully considered the potential risks to both the participants of the survey.
Risk Mitigation.The main risk for the survey participants is the breach of privacy. To mitigate this risk, we ensured that no identifiable data (e.g., names, contact details) was collected from survey participants, and the survey responses remain fully anonymous. Additionally, the survey itself includes an informed consent section that informs the respondents about the voluntary participation, the survey’s purpose, a description of the procedures, the risks involved, contact information, and the option to opt out of the survey. Finally, we provided participants the choice to skip questions using the response options “Prefer not to say” and “I do not know” to applicable questions. Due to the aforementioned risk mitigation strategies, the stakeholders’ breach of privacy risk is negligible.
@@ IX-B Honeypot Deployment Experiment
Stakeholder Identification. The primary stakeholders in the honeypot deployment are the external actors who interact with our honeypot and the infrastructure owners on whose infrastructure our honeypot is running. In the honeypot deployment experiment we deployed HoneySat on server infrastructure owned by real-world institutions.
Risk Mitigation. The main risk for the infrastructure owners is that one of the external actors will use our honeypot as an stepping stone to breach their production infrastructure. For the deployment on the premises of the institutions, we worked with the local IT infrastructure administrators to ensure that we took all the necessary precautions to avoid any possibility that the external actors could gain access to the infrastructure via our honeypot. First, the infrastructure owners provided us with a server with a clean OS installation that did not have any production applications or sensitive data. Second, the server was isolated on its own network segment. Third, we configured a firewall to only allow the necessary ports. Fourth, the VNC portion of the honeypot is configured so that external actors cannot use it to host malicious services by denying traffic forwarding. Fifth and final, we deployed our honeypot using two sandboxing layers (Docker containers and Virtual Machines (VMs) as we discussed in Appendix A-C. Due to the aforementioned risk mitigation strategies, the stakeholders’ breach of privacy risk is negligible.
The main risk for the external actors is the breach of their privacy. However, it is generally accepted that trespassers of a computer system do not have reasonable expectation of privacy [78]. In order to mitigate this risk, our honeypot system gives notice and warning to any user connecting to it indicating that ”This computer system is for authorized use only.” Additionally, the web services of our honeypot are protected by a username and password. This notice system gives external actors a chance to avoid interaction, which helps reduce the risk.
Appendix A Appendix
A-A Interaction Sequences and Exploits
@@
Exploit OS, libraries or software vulnarabilities.
@@
tm_send_file articaft 1
@@
Scape the ground software or docker and run tcpdump.
@@
Scape the ground software or docker and run tcpdump.
@@
Deploy an malicious application that sends data over CSP
Command and Control Telecommand a Spacescraft T2019
@@
The atacker deploys a custom CSP node or a backdoor
@@
Send TC to change communication paramters.
@@
Send TC to lauch a fork bomb or a reset loop
@@
A-B Satellite Simulator Design Details
Figure 6: High-level architecture of the Satellite Simulator. Due to space limitations, we do not show all the available sensors and simulations. Some sensors communicate with multiple simulations. Although the flight software services are not part of the Satellite Simulator we included it to depict the Interface module’s data flow better.
As shown in Fig. 6, the Satellite Simulator includes five module types: simulation modules (light orange), sensor modules (light blue), subsystem state modules (light green), satellite state modules (light gray) and interface modules (light yellow).
Simulation Modules. These modules are abstractions of real-world processes required for a realistic satellite simulation (DP-2). An example of these simulations is the orbital simulation. This simulation uses orbital mechanics to calculate data such as the satellite’s orbital position at any given time.
These simulation modules can communicate with each other to facilitate proper functionality. For example, the power simulation module will query the orbital simulation for orbital data to determine if the satellite is positioned properly in relation to the sun so that it can draw power from its solar panels. The subsystems and the sensors of the Satellite Simulator are designed around these simulations. The Satellite Simulator includes the orbital, rotation, power system, thermal, magnetic, and payload simulations. Due to space restrictions and their complex design, we explain each simulation in Appendix A-B.
Overall, these simulations provide the foundation for any satellite operator to use our framework and implement their honeypot using their satellite flight software. The Satellite Simulator simulations are configured using the satellite personality, which we discuss below.
To finish, the Satellite Simulator’s built-in modular design allows subsystems, simulations, and sensor modules to be mixed ’n matched to meet the requirements of most satellites and their missions. Additionally, our flexible design allows users to easily add additional simulations, sensors, or subsystems as modules to the honeypot.
  • Orbital Simulation. This simulation calculates and predicts the satellite’s orbit. To achieve this, it uses the two-line element set or TLE data configured in the satellite personality. The TLE data contains the necessary orbital information required to calculate the location of a satellite given a point in time [43]. We rely on the Skyfield python library to perform advanced and precise calculations [79]. This simulation calculates multiple orbital values which include the altitude angle, latitude, longitude of the satellite among other values. The payload, EPS, and magnetic simulations rely on data from this simulation.
  • Rotation Simulation. This simulation simulates the satellite attitude changes to provide the satellite orientation. Determining and controlling the satellite orientation is relevant for payloads such as cameras or communication systems that require pointing to the earth surface or to other celestial bodies. This simulation calculates the rotation kinematics of the satellite using two reference frames: a reference frame fixed to the satellite body and a non-rotational reference frame. The relation between the two reference frames is calculated using the conservation of angular momentum and the rigid-body Euler equation [80]. To simulate the satellite’s rotation, the initial instantaneous angular velocity of the satellite, the current orientation in quaternions, and the satellite’s inertia matrix must be defined. The simulation uses the fourth-order Runge Kutta integration method (RK4) [81] to perform these calculations to avoid divergences in the simulation with time steps greater than 1 second.
  • Power System Simulation. This simulation manages the satellite’s power collection, consumption, and distribution. It tracks the battery capacity, the power draw and simulates battery charging whenever the orbital simulation tells it that the satellite is exposed to sunlight. A battery pack is modeled as a series or parallel configuration of smaller battery packs. The simplest battery pack consists of a single cell. A single cell is simulated to calculate the charge and discharge curves, assuming that the cell’s current will not change. This cell voltage is then used to calculate the pack’s voltage. The electrical power system simulation uses the pack’s voltage together with the input and output power to calculate the battery current. In case the input or output power of the power subsystem changes, a new cell current is calculated, and the battery cell simulation is rerun to reflect the change.
  • Thermal Simulation. This simulation calculates the temperature of the satellite. The simulation is based on two values: the total thermal energy and a user-defined specific heat capacity. Together with emissivity, mass and surface area, the temperature of a satellite can be approximated based on the absorbed sunlight and the energy emitted through thermal radiation. This simulation uses a radiation loss formula [82] to calculate the thermal radiation emission.
  • Magnetic Simulation. This simulation tracks and analyzes the interactions between the Earth’s magnetic field and the satellite’s own magnetic environment. It communicates with the orbital and rotation simulations to determine the satellite’s position and orientation in space so that it can calculate the Earth’s magnetic field components (north, east, vertical) and total intensity at the satellite’s current location.
  • RGB Camera Simulation. This simulation replicates the functionality of an RGB camera pointed at Earth. It uses Earth observation satellite imagery from the EarthExplorer database from the U.S. Geological Survey (USGS) [77]. If a capture image command is issued, it will select an image from one of two pools of images: daytime or nighttime. This simulation communicates with the orbital simulation to determine if satellite is in the presence of the sun.
A-C Deployment and Security Hardening Implementation
Figure 7: Architecture of dockerized Generic CSP Honeypot
Containerization. After we completed HoneySat’s development, we used Docker Compose [83] to containerize each of our framework’s applications. Specifically, we created four different containers depicted in Fig. 7. Containerizing HoneySat provides two benefits. First, it creates another sandboxing layer that prevents adversaries from using our honeypot to access the underlying system [32]. Second, it proves a convenient and flexible way to deploy HoneySat.
Figure 8: Screenshot of the HoneySat YAMCS web interface, displaying a TM packet, returned from the Generic CCSDS Mission Honeypot.
A-D Extensibility Example Use Cases
TABLE IX: HoneySat example use cases for further extensibility.
Extensibility Use Case Relevant HoneySat Module(s) High-level Steps
Add new sensor Satellite Simulator To add a new sensor you can leverage the existing Sensor Abstract Base Class (ABC) included in the Satellite Simulator source code and write the logic for your new sensor.
Add new payload Satellite Simulator To add a new payload you can leverage the existing PayloadSimulation Abstract Base Class (ABC) included in the Satellite Simulator source code and write the logic for your new payload including where to get the payload data.
Integrate new flight software Satellite Simulator, ground configuration, satellite personality, radio simulator, flight software services To integrate a new flight software into HoneySat the first step is to analyze and understand the target flight software architecture and source code even before attempting integration. Then the target flight software must be integrated using HoneySat’s flight software services module. After integrating the flight software services, the Satellite Simulator can be connected using the Interface Python ABC locanted in the Satellite Simulator source code.
Add a new exposed network protocol Ground Software, Docker compose A new exposed network protocol can be added by modifying the Docker compose file included with HoneySat. The compose file allows for new services and TCP ports to be easily added.

### v3

@@ VII-A Initial Satellite Honeypot Anti-Detection Techniques
Limitations. Currently, the functionality of some subsystems within HoneySat’s Satellite Simulator are constrained by the quality of the data provided. For example, the resolution of Earth’s images generated by our camera payload depends on the resolution of its source data (USGS [geological2022earthexplorer]). Consequently, creating a honeypot for a satellite with a high-resolution camera payload would not be feasible without addressing the underlying issues of data source quality.
@@ VIII Conclusion
Although we have yet to witness a Stuxnet-like cyberattack on space systems, security researchers need to develop effective countermeasures to secure satellites. In this paper, we introduced HoneySat, the first satellite honeypot that simulates a satellite mission and provided evidence that our honeypot can obtain real-world interaction data and can be extended with real-world satellite software and hardware. Finally, we hope that security researchers use HoneySat’s open-source implementation as a foundation not only for satellite honeypot deployments but also for simulation, and training applications.
Acknowledgments
We thank our shepherd and the anonymous reviewers for their helpful suggestions towards improving this paper. The project underlying this paper was funded by the Federal Ministry of Transport (BMV) under the code 45AVF5A011 and was partially supported by NSF awards No. 2131263 and 2232911, and by the US Department of Transportation (USDOT) CYBER-CARE Grant No. 69A3552348332. We further thank the Saarbrücken Graduate School of Computer Science for their funding and support. The authors are responsible for the content of this publication.
@@ IX Ethical Considerations
In this paper, we consider the ethical consequences and possible negative outcomes of the satellite operator user study and the HoneySat deployments. We now discuss the stakeholders, potential risks, and how we mitigated those risks.
@@ IX-A Satellite Operators User Study
Prior to conducting any research, the survey protocol, and materials were submitted for review by our Institutional Review Board (IRB). The protocol received “exempt status”, indicating no more than minimal risk.
Stakeholder Identification. The primary stakeholders in the survey are survey participants.
Risk Mitigation. The main risk for the survey participants is the breach of privacy. To mitigate this risk, we ensured that no identifiable data (e.g., names) was collected, and the responses remain anonymous. Additionally, the survey itself includes an informed consent section that informs the respondents about the voluntary participation, the survey’s purpose, a description of the procedures, the risks involved, contact information, and the option to opt out of the survey. Finally, we provided participants the choice to skip questions using the response options “Prefer not to say” and “I do not know.”
@@ IX-B Honeypot Deployment Experiment
Stakeholder Identification. The primary stakeholders in the honeypot deployment are the external actors who interact with our honeypot and the infrastructure owners on whose infrastructure our honeypot is running.
Risk Mitigation. The main risk for the infrastructure owners is that external actors will use our honeypot as a stepping stone to breach their infrastructure. For the on-prem deployments, we worked with the local IT administrators to ensure that we took all the necessary precautions to avoid this scenario. First, we were provided with a server with a clean OS installation that did not have any production applications or sensitive data. Second, the server was isolated on its own network segment. Third, we configured a firewall to only allow the necessary ports. Fourth, the VNC portion of the honeypot is configured so that external actors cannot use it to host malicious services by denying traffic forwarding. Fifth and final, we deployed our honeypot using two sandboxing layers (Docker containers and Virtual Machines (VMs) as we discussed in Appendix A.
The main risk for the external actors is the breach of their privacy. However, it is generally accepted that trespassers of a computer system do not have reasonable expectation of privacy [spitzner2002honeypots]. In order to mitigate this risk, our honeypot system gives notice and warning to any user connecting to it indicating that “This computer system is for authorized use only.” Additionally, the web services of our honeypot are protected by a username and password.
Appendix A Deployment and Security Hardening Implementation
Figure 6: Architecture of dockerized Generic CSP Honeypot
Containerization. After we completed HoneySat’s development, we used Docker Compose [docker2024docker] to containerize each of our framework’s applications. Specifically, we created four different containers depicted in Fig. 6. Containerizing HoneySat provides two benefits. First, it creates another sandboxing layer that prevents adversaries from using our honeypot to access the underlying system [marra2024feasibility]. Second, it proves a convenient and flexible way to deploy HoneySat.
TABLE VII: Correspondence between real satellite mission components and HoneySat’s design components.
Segment Real Component HoneySat Component
Ground Remote Operation Protocols VNC, Telnet
Ground Mission Control Software Modified MCS
Ground Ground Station Control Software Gpredict
Ground Web Interface Web Interface
Ground Ground Station Radio Simulator
Space Flight Software F.S. Runtime + Services
Space On Board Computer Docker Container
Space Platform Satellite Simulator
Space Payload Satellite Simulator
Appendix B Interaction Sequences and Exploits
@@
Exploit OS, libraries or software vulnerabilities.
@@
tm_send_file artifact 1
@@
Escape the ground software or docker and run tcpdump.
@@
Escape the ground software or docker and run tcpdump.
@@
Deploy a malicious application that sends data over CSP
Command and Control Telecommand a Spacecraft T2019
@@
The attacker deploys a custom CSP node or a backdoor
@@
Send TC to change communication parameters.
@@
Send TC to launch a fork bomb or a reset loop
@@
Appendix C Hardware-in-the-loop Experiment Results
Fig. 7 depicts a snippet of radio signal data which confirms that the hardware-in-the-loop experiment in Sec VI-F was successful. This was provided by the aerospace company.
Appendix D Artifact Appendix
This appendix accompanies the paper “HoneySat: A Network-based Satellite Honeypot Framework” and provides detailed instructions for obtaining, installing, and evaluating the artifact submitted for NDSS 2026 Artifact Evaluation.
D-A Description & Requirements
Our artifact contains a Dockerized version of our honeypot framework, along with all the necessary components to bootstrap honeypots for satellite missions supporting two different space protocol stacks (CSP and CCSDS/YAMCS). To aid in evaluations, we also included utility scripts that facilitate the evaluation of data provided by the honeypot.
D-A1 How to access
The artifact can be found at https://doi.org/10.5281/zenodo.17871431.
D-A2 Hardware dependencies
  • 25GB of disk space
  • at least 4 CPU cores
  • 8GB RAM
D-A3 Software dependencies
Most Linux distributions will work. We verified the artifact on Ubuntu 24.04.2 LTS.
  • Docker (with compose and related components, we recommend to follow https://docs.docker.com/engine/install/)
  • Python 3.12 (more recent versions may also be compatible)
  • Python virtual environment
  • Bash
  • Telnet client
  • Any modern web browser
D-A4 Benchmarks
No External Benchmarks required.
D-B Artifact Installation & Configuration
Extract files
Extract files from tar file downloaded from link:
tar xzpvf ndss-artifact-eval.tar.gz
Repository Layout
The root directory contains two subdirectories:
  • deployment/: Dockerized HoneySat services with detailed setup documentation.
  • evaluation/: Python utilities and scripts to help evaluate HoneySat’s capabilities.
The evaluation utilities require Python 3.12 and the packages listed in evaluation/requirements.txt.
Create and Activate a Python Virtual Environment.
From the repository root:
python3 -m venv .
source bin/activate
Install Python Dependencies
Install the required packages for the evaluation utilities:
python3 -m pip install -r evaluation/requirements.txt
Configure Docker user
We need to add the current user to the Docker group to avoid using sudo with Docker all the time.
sudo usermod -aG docker \$USER
sudo reboot now
Start the Honeypot Services.
All remaining setup is handled by Docker containers. You can either:
  • Run the provided convenience scripts (see “Quick Start for Pass Simulations” below), or
  • Follow the guide in deployment/README.md to bring up the services via Docker Compose.
Quick Start for Pass Simulations
A key HoneySat feature is its realistic communication windows: the satellite is reachable only while passing over a ground station. Many experiments benefit from a setup in which the satellite becomes reachable immediately or within a short time window. We provide two helper scripts that compute a suitable ground-station location along the satellite’s predicted ground track so that a pass occurs shortly after startup. Each script builds and starts the necessary Docker Compose services and will shut them down cleanly on CTRL+C.
  • CSP-based honeypot:
    ./evaluation/experiment-1/run-experiment-csp.sh
    This script builds and starts a CSP honeypot instance, and positions the ground station to enable communication almost immediately, allowing you to observe how an attacker would perceive a pass without waiting for a real one.
  • CCSDS/YAMCS-based honeypot:
    ./evaluation/experiment-1/run-experiment-ccsds.sh
    This script builds and starts a CCSDS and YAMCS-based honeypot instance, using the same predicted ground-station placement to trigger an imminent pass.
Please proceed with the experiments described in the following sections, or consult deployment/README.md for additional configuration details.
D-C Experiment Workflow
The following experiments all involve starting the honeypot system and interacting with it in some way (either programmatically or manually). The workflow typically involves issuing a startup command and then performing tasks outlined in the experiment description. For some tasks, timing matters; for example, the TCs have to be issued while the simulated satellite is reachable via the simulated ground station.
D-D Major Claims
We make the following claims in our paper
  • (C1): HoneySat can be used to deceive adversaries and log activities.
    • HoneySat’s simulator provides believable TM (Experiment 1.1)
    • HoneySat has realistic communication windows (Experiment 1.2)
    • HoneySat provides interaction capabilities (process TC, provide TM) (experiment 1.3)
    • HoneySat Logs interaction details (Experiment 1.4)
  • (C2): HoneySat Is extensible and supports two different protocol ecosystems
    • HoneySat is configurable (Experiment 2)
D-E Evaluation
D-E1 Experiment 1 (E1.1)
[Believable TM/TC]   [10 human-minutes]
Goal. Demonstrate that HoneySat produces believable telemetry (TM) in response to telecommands (TC) for a CCSDS/YAMCS setup.
Preparation. Ensure dependencies are installed as listed above.
Execution.
./evaluation/experiment-1/run-experiment-ccsds.sh
After the script completes and prints telemetry values, you may also inspect the telemetry via the YAMCS web interface.
Expected Results. Believable current, voltage, and temperature values for the selected battery configuration:
  • Voltage:  8000 mV (normal test case)
  • Temperature:  30 °C
  • Current draw:  74 mA
D-E2 Experiment 1.2 (E1.2)
[Believable passes]   [20 human-minutes]
Goal. Show that HoneySat enforces realistic communication windows: the satellite is reachable only during predicted passes over a ground station.
Preparation.
./evaluation/experiment-1/run-experiment-csp.sh
This positions the ground station so that a pass begins approximately 2-3 minutes after startup (adjustable), and starts all required services.
Execution.
  1. 1.
    Open the web interface at http://localhost:80. Log in with username admin and password admin. Click the ground-station icon and view the next predicted passes. The pass should soon show as “ongoing.”
  2. 2.
    In a separate terminal, connect to the CSP telnet interface and activate it:
    telnet localhost 24
    # In the telnet session:
    activate
  3. 3.
    Probe satellite reachability every few seconds:
    1: com_ping 10
Expected Results. The satellite responds to pings only during the predicted pass and not before or after. Responses will indicate the expected addressing (source address 1, destination address 10).
D-E3 Experiment 1.3 (E1.3)
[Simulate Interaction]   [15 human-minutes]
Goal. Demonstrate interactive capabilities once the satellite becomes reachable.
Preparation. Repeat the setup from Experiment 1.2. Review the command reference in the experiment-1.3 directory.
Execution. After the satellite becomes reachable (as verified in Experiment 1.2), issue commands via the telnet interface. For example, to execute arbitrary shell commands on the OBC:
1: obc_system [shell command]
Expected Results. The experimenter can successfully execute arbitrary shell commands via obc_system while the satellite is reachable.
D-E4 Experiment 1.4 (E1.4)
[View logging capabilities]   [5 human-minutes]
Goal. Verify that HoneySat logs interactions and relevant parameters.
Preparation. Keep a HoneySat CSP instance running (e.g., from Experiment 1.3), optionally after interacting with it.
Execution.
python3 ./evaluation/experiment-1/python_dump_mongodb/dump_mongodb.py
Alternatively, connect to the MongoDB instance with a database client of your choice.
Expected Results. The script prints MongoDB contents to stdout, showing logged parameters and interaction details.
D-E5 Experiment 2 (E2)
[Customization]
Goal. Demonstrate HoneySat’s configurability across protocol stacks and scenarios.
Preparation. Navigate to ./evaluation/experiment-2.
Execution.
  1. 1.
    Create a baseline customization:
    python3 ./honeysat.py [csp|ccsds] "{satellitename}" "{location}"
    For example:
    python3 ./honeysat.py csp "BEESAT" "Berlin"
  2. 2.
    Start services with the generated configuration (or configs in this directory):
    python3 ./honeysat.py start [csp|ccsds] "{satellitename}" "{location}"
    For example:
    python3 ./honeysat.py start csp "BEESAT" "Berlin"
  3. 3.
    When finished, stop the services:
    python3 ./honeysat.py stop [csp|ccsds]
    For example:
    python3 ./honeysat.py stop csp
Expected Results. HoneySat can be configured with relatively low time effort to reflect different satellites, locations, and protocol ecosystems.